# MDCure: A Scalable Pipeline for Multi-Document Instruction-Following

### v1

@@ 4 Experimental Setup
  • The scalability of MDCure as we vary the size of the generated instruction data and the size of the base LLM being instruction-tuned.
@@
  • The importance of the Filtering stage for curating high-quality MD instruction data.
  • The superiority of MDCure’d models versus pre-trained MD and long-context baselines in 0-shot prompting settings.
@@
  • The benefits of MDCure over existing synthetic MD-focused pre-training datasets.
@@
  • The generalizability of MDCure to few-shot and long-context MD instruction data generation.
  • The value of different prompt approaches toward generating effective MD instructions.
In addition, we conduct ablations to demonstrate the criticality of each of the MDCureRM scoring criteria and the effect of their relative weightings. We discuss alternative architectures and the design of MDCureRM in App. B.2. We further provide a meta-evaluation to analyze the alignment of MDCureRM with human annotations versus GPT-3.5-Turbo in App. D, where we observe moderate agreements in line with LLM evaluation literature (Liu et al., 2024b). As we discuss later, we observe strong indication that MDCureRM provides consistent empirical improvements.
@@
Models and Training Details. We apply MDCure on several LLMs across different widely used model families and sizes: FlanT5 (Chung et al., 2022), Qwen2 (Yang et al., 2024), and LLAMA3.1 (Dubey et al., 2024). We chose these models as they are capable open-source instruction-following models with varying architectural designs and parameter scales. We utilize the instruction-tuned variants as our base models to showcase the complementary effect of MDCure instructions versus typical IT data. To assess the performance of MDCure at scale, we create MDCure instruction datasets of size 12000, 36000, and 72000444Dataset sizes were determined as a function of API costs and based on the sizes of representative IT datasets used in related works (Bai et al., 2024; Köksal et al., 2023). and use them to fine-tune FlanT5-Base and -Large (250M & 750M), Qwen2-Instruct (1.5B & 7B), and LLAMA3.1-Instruct (8B & 70B). Hereafter, we refer to resulting models that have undergone MDCure IT as “MDCure’d” models. To assess the efficacy of MDCureRM, we additionally repeat these experiments using unfiltered and GPT-3.5-filtered instruction sets. [^5] We utilize GPT-3.5-Turbo for comparison as it is a more cost-effective alternative to commonly used scorer GPT4. Cheaper models such as GPT-4o-mini had not yet been released at the time of our starting experimentation. Further details regarding data and training setup are in App.s A and B.1.
@@
Baselines. Beyond their base model counterparts, our MDCure’d models are compared against state-of-the-art pre-trained models for multi-document processing, namely PRIMERA and QAMDen, which are comparable in scale to FlanT5-Base and -Large, as well as several strong open-source long-context LLMs, namely LongAlign-7B (Bai et al., 2024) and ProLong-8B-64k-Instruct (Gao et al., 2024b), which are comparable to Qwen2-7B-Instruct and LLAMA3.1-8B-Instruct, and Jamba-1.5B-Mini (Labs, 2024) which has 54B parameters. We also compare MDCure’d models to SOTA general LLMs GPT-4o and Gemini 1.5 Pro (Team et al., 2024). Additionally, we inspect the effect of our instruction data versus baseline multi-document and long-context pre-training data at the 7B- and 8B-scales by fine-tuning Qwen2-7B-Instruct with PRIMERA and QAMDen data and ProLong-8B-64k-Instruct with MDCure data (72k). Further details regarding baselines are in App. C.2.
@@
Benchmarks. We evaluate all baselines and MDCure’d models over a range of long- and short-form tasks through use of two challenging MD and long-context benchmarks: SEAM (Lior et al., 2024) and ZeroScrolls (Shaham et al., 2023). Within SEAM, we specifically consider the MultiNews, OpenAsp, MuSiQue, ECB+, and SciCo datasets to ensure coverage of a spread of domains and MD tasks. We additionally profile 0-shot prompting performance over a suite of four widely used multi-document benchmarks: WikiHop (Welbl et al., 2018) and HotpotQA-distractor (Yang et al., 2018) for MHQA, Multi-XScience (Lu et al., 2020) for MDS, and QMDSCNN (Pasunuru et al., 2021) for QMDS. While these datasets were built for use in fine-tuning evaluation, we draw upon the prompting approach proposed by Shaham et al. (2023) to formulate effective 0-shot prompts for each model. App.s C.1 and C.3 provide additional details.
@@
Metrics. For SEAM and ZeroScrolls we employ the metrics and corresponding implementations defined by each for their component datasets. For WikiHop and HotpotQA-distractor, we use the standard F1 score. For Multi-XScience and QMDSCNN, we utilize LLM scores issued by GPT-3.5-Turbo and Gemini 1.5 Flash, respectively, to assess summarization coherence, relevance, fluency, and consistency. This follows recent adoption of LLM-as-a-Judge as a proxy for human-level rating (Liu et al., 2024c; Ye et al., 2024), given the limitations of traditional metrics like ROUGE for long-form generation quality assessment. (Sai et al., 2019; Gao et al., 2024a). Further details regarding the evaluator LLMs and the prompts are in App. C.4.
@@ 5 Results
IT Data Setting HQA WH MXS QC SEAM ZS Score Avg
@@
FlanT5 Base None 4.4 45.1 1.74 2.16 0.0 5.8 0.2 2.6 0.1 1.7 13.1 10.2
Unfiltered 31.7 47.5 4.04 2.35 0.0 6.7 0.1 2.7 0.2 1.9 21.0 16.1
GPT-Filtered 44.1 46.0 4.16 2.44 0.0 6.0 0.2 2.7 0.0 1.8 21.3 16.8
MDCureRM 47.3 48.3 4.22 2.58 0.0 7.5 0.3 2.7 0.0 2.1 22.6 17.9
Large None 24.4 54.6 3.19 2.78 1.1 7.9 0.1 3.1 0.3 2.5 23.2 17.3
Unfiltered 46.9 62.9 4.10 2.91 0.7 8.0 0.0 3.8 0.1 2.5 24.2 19.5
GPT-Filtered 46.3 65.1 4.13 2.88 0.5 8.3 0.0 3.9 0.1 2.6 24.2 19.7
MDCureRM 49.6 66.1 4.19 2.97 1.2 9.1 0.1 4.2 0.0 2.9 25.3 20.5
Qwen2-Ins 1.5B None 21.5 17.8 4.20 3.30 9.4 10.6 0.5 5.0 13.0 7.7 24.0 17.1
Unfiltered 32.9 30.5 4.24 3.57 15.7 12.0 0.4 5.0 16.8 10.0 23.9 18.9
GPT-Filtered 33.3 32.9 4.24 3.66 16.1 12.0 0.5 5.1 15.9 9.9 24.2 19.3
MDCureRM 37.7 34.8 4.25 3.73 16.4 12.0 0.6 5.9 18.1 10.6 25.6 20.5
7B None 30.5 39.6 4.30 3.57 5.0 11.9 0.5 6.4 13.1 7.4 23.9 18.6
Unfiltered 40.5 43.3 4.26 3.79 8.0 15.3 0.5 6.6 11.9 8.5 26.3 20.9
GPT-Filtered 42.0 44.0 4.26 3.84 8.7 15.3 0.9 6.6 12.1 8.7 28.7 22.3
MDCureRM 44.7 46.0 4.28 3.93 13.8 15.4 0.6 6.7 14.9 10.3 29.8 23.6
+PRIMERA 44.6 45.4 3.96 2.17 5.9 15.3 0.5 6.6 12.4 8.1 27.7 21.8
+QAMDen 42.6 45.8 3.92 2.37 9.5 15.3 0.5 6.4 12.4 8.8 28.5 22.3
Llama3.1-Ins 8B None 35.5 27.1 4.28 2.94 10.5 15.0 0.6 7.6 17.4 10.2 18.7 16.2
Unfiltered 37.6 34.4 3.81 4.07 15.3 16.3 0.5 7.8 18.9 11.8 28.6 22.3
GPT-Filtered 38.0 42.3 4.29 3.95 8.7 16.0 0.6 7.8 18.3 10.3 29.6 22.9
MDCureRM 44.7 43.7 4.29 4.22 16.3 16.4 0.6 7.9 18.5 11.9 30.9 24.5
70B None 53.9 38.1 4.28 3.97 25.1 21.9 0.6 6.1 11.5 13.0 36.4 27.9
Unfiltered 55.9 40.6 3.99 3.17 3.6 21.9 0.6 6.3 11.2 8.7 37.2 27.3
GPT-Filtered 57.4 41.2 3.97 3.37 4.9 22.0 0.7 6.4 10.7 8.9 37.4 27.6
MDCureRM 58.4 45.5 4.28 3.99 25.2 22.1 0.7 6.7 12.0 13.3 37.7 29.2
    Baselines PRIMERA 0.4 0.5 3.18 1.09 7.1 7.7 0.5 3.9 10.4 5.9 4.2 4.1
QAMDen 1.8 1.9 2.86 1.22 0.0 0.4 0.0 1.0 3.1 0.9 3.8 2.6
LongAlign-7B 10.4 14.3 4.15 3.75 11.5 16.5 0.0 4.1 16.8 9.8 23.2 16.5
ProLong-8B 43.6 34.3 3.86 2.46 9.5 17.4 0.4 10.7 18.0 11.2 32.2 24.3
Jamba 1.5 Mini 47.5 41.8 4.24 3.92 20.1 14.2 0.0 5.3 20.4 12.0 35.1 26.8
GPT-4o 57.5 50.0 4.50 4.25 8.9 18.6 0.3 14.2 28.9 14.1 39.8 30.8
Gemini 1.5 Pro 66.6 55.6 4.22 3.59 21.4 17.8 0.6 15.1 30.1 17.0 32.0 28.2
Table 1: Evaluation of MDCure’d models versus baselines on 6 benchmarks in the zero-shot prompting setting. Dataset abbreviations are described in App. C.1.1. Full results on ZeroScrolls are provided in App. E. Rows specified by MDCureRM refer to our full MDCure pipeline applied to the corresponding base model and size. Bold numbers show best performance in each group.
@@ 5.1 Main Results
We report experimental results in Table 1, with full results on ZeroScrolls included in App. E. For fair comparison, we report all MDCure results based on training with 72K samples, which led to best performance overall across all models. Our key findings are summarized as follows.
MDCure instructions are effective across model families and sizes. MDCure exhibits the best performance across all base models, demonstrating a clear and consistent advantage relative to the non-MDCure’d base models over all benchmarks. In particular, we see improvements of up to 975%, 96%, 143%, 434%, 39%, and 73% for HotpotQA, WikiHop, Multi-XScience, QMDSCNN, SEAM, and ZeroScrolls, respectively. MDCure’d LLAMA3.1-70B-Instruct achieves the best results across the 6 benchmarks, followed by MDCure’d Qwen2-7B-Instruct. As our models are trained from instruct-versions of base LLMs, it is clear that MDCure instruction data imparts complementary abilities versus typical IT data.
The relative improvement of MDCure’d models versus corresponding base models varies across model families, and improvements tend to shrink with model size. For FlanT5, MDCure provides 75.5% average improvement at the 250M-parameter scale (FlanT5-Base) versus 18.5% at the 750M-parameter scale (FlanT5-Large). For LLAMA3.1, MDCure provides 51.2% average improvement at the 8B scale versus 4.7% at the 70B scale. For Qwen2, we observe alternatively that MDCure gains shift from 19.9% average improvement at the 1.5B scale to 26.9% at the 7B scale. More generally, these results suggest the importance of instruction quality may weaken with base model scale and capability, similar to findings by Ivison et al. (2023).
Filtering with MDCureRM is essential to improve MD capabilities of LLMs. Comparing to models trained with unfiltered MDCure instruction samples, MDCure’d models which have been trained with the highest-scoring MDCureRM-filtered instructions surpass them across all benchmarks and base models. Furthermore, MDCureRM outperforms GPT-3.5-Turbo as a judge of MD instruction quality and efficacy, evidenced by the superior performance of MDCure’d models versus those trained with GPT-filtered instructions.
MDCure improves LLM performance on both multi-document tasks and single-document long-context tasks. Across MD benchmarks, MDCure enables models to achieve improved performance in all settings. In addition to MD tasks, MDCure generally boosts single-document long-context performance, particularly in tasks from ZeroScrolls, which emphasize general long-context capabilities. These trends are consistent across downstream task domains (e.g., science, literature), suggesting the MDCure procedure can boost MD performance in a domain-general way.
@@
ProLong (PL) 43.6 34.3 3.86 2.46 11.2 32.2 24.3
PL+MDCure 45.8 45.1 3.93 3.40 12.5 32.2 25.4
Table 2: Our MDCure pipeline’s performance when applied to an already strong long-context LLM, ProLong-8B (Gao et al., 2024a).
Our best MDCure’d models consistently outperform pre-trained long-context baselines. For all benchmarks, our MDCure’d models achieve results that exceed the performance of size-comparable baselines [^6] We consider Jamba 1.5 Mini (54B) comparable with LLAMA3.1-70B-Instruct, as their parameter scales are within the same ballpark, falling within the same order of magnitude. by a substantial margin, confirming the efficacy of our design. For a more controlled setup, we take a strong open-source long-context LLM Prolong-8B (Gao et al., 2024a)) and fine-tune it using MDCure. The goal of this experiment is to study if our data provides additional advantages to long-context continual pre-training and fine-tuning. Table 2 demonstrates the results. Note that Prolong has been trained on large datasets totaling $\sim$81B tokens to improve long-context performance. As we observe, fine-tuning ProLong-8B with MDCure data which is much smaller ($\sim$162M tokens) yields further improved performance, indicating the complementary value of our MD training data to long-context training.
Our MDCure pipeline improves over existing multi-document synthetic data pipelines. We also compare our MDCure pipeline with existing MD-focused synthetic pre-training datasets inluding PRIMERA (Xiao et al., 2022) and QAMDen (Caciularu et al., 2023). As indicated by the “+PRIMERA” and “+QAMDen” rows in Table 1, MDCure data offers performance improvements across benchmarks at the 7B-scale. The weak performance of the original PRIMERA and QAMDen models (in the Baselines rows) can be explained as they are intended to serve as strong base pre-trained models for finetuning as opposed to zero-shot prompting settings. Overall, our results emphasize the utility of synthetic MD instructions as a promising approach for improving MD capabilities of LLMs.

### v3

@@ 4 Experimental Setup
  • The scalability of MDCure as we vary the size of the generated instruction data and the size of the base LLM being post-trained.
@@
  • The importance of MDCureRM filtering for curating high-quality MD instruction data.
  • The strong cross-task and cross-domain generalization and superior results achieved by MDCure over pre-trained and long-context baselines.
@@
  • The ability for MDCure data to excel over existing synthetic MD pre-training datasets.
@@
  • The adaptability of MDCure for few-shot and extended-context MD instruction data generation.
  • The compatibility and efficacy of MDCure with open- and closed-source generator LLMs.
  • The value of MDCureRM toward providing reward signals for synthetic MD data policy optimization.
@@
Models and Training Details. We apply MDCure to LLMs across different widely used model families and sizes: FlanT5, Qwen2, and LLAMA3.1. We choose these models as they are capable open-source instruction-following models with varying architectures and parameter scales. We utilize the instruction-tuned variants as our base models to showcase the complementary effect of MDCure instructions versus typical IT data. To assess MDCure at scale, we create MDCure instruction datasets of size 12000, 36000, and 72000 examples [^5] Dataset sizes were determined as a function of API costs and based on the sizes of representative IT datasets used in related works (Bai et al., 2024; Köksal et al., 2023). and use them to fine-tune FlanT5-Base and -Large (250M & 750M), Qwen2-Instruct (1.5B & 7B), and LLAMA3.1-Instruct (8B & 70B). Hereafter, we refer to resulting models that have undergone MDCure post-training as “MDCure’d” models. To assess the efficacy of MDCureRM, we additionally repeat these experiments using unfiltered and GPT-3.5-filtered instruction sets. [^6] We utilize GPT-3.5-Turbo for comparison as it is a more cost-effective alternative to commonly used scorer GPT4. Cheaper models such as GPT-4o-mini had not yet been released at the time of our starting experimentation. Further data and training setup details are in App.s B and C.1.
@@
Baselines. Beyond their base model counterparts, our MDCure’d models are compared against state-of-the-art pre-trained models for multi-document processing, namely PRIMERA (Xiao et al., 2022) and QAMDen (Caciularu et al., 2023), which are comparable in scale to FlanT5-Base and -Large, as well as several strong open-source long-context LLMs, namely LongAlign-7B (Bai et al., 2024) and ProLong-8B-64k-Instruct (Gao et al., 2024b), which are comparable to Qwen2-7B-Instruct and LLAMA3.1-8B-Instruct, and Jamba-1.5B-Mini (Labs, 2024) which has 54B parameters. We also compare MDCure’d models to SOTA general LLMs GPT-4o and Gemini 1.5 Pro (Team et al., 2024). Additionally, we inspect the effect of our instruction data versus baseline MD and long-context pre-training data at the 7B- and 8B-scales by fine-tuning Qwen2-7B-Instruct with PRIMERA and QAMDen data and ProLong-8B-64k-Instruct with MDCure data (72k). Further details regarding baselines are in App. D.2.
@@
Benchmarks. We evaluate all baselines and MDCure’d models over a range of long- and short-form tasks through use of two challenging MD and long-context benchmarks: SEAM (Lior et al., 2024) and ZeroScrolls (Shaham et al., 2023). Within SEAM, we specifically consider the MultiNews, OpenAsp, MuSiQue, ECB+, and SciCo datasets to ensure coverage of a spread of domains and MD tasks. We additionally profile 0-shot prompting performance over a suite of four widely used MD benchmarks: WikiHop (Welbl et al., 2018) and HotpotQA-distractor (Yang et al., 2018) for MHQA, Multi-XScience (Lu et al., 2020) for MDS, and QMDSCNN (Pasunuru et al., 2021) for QMDS. While these datasets were built for use in fine-tuning evaluation, we draw upon the prompting approach proposed by Shaham et al. (2023) to formulate effective 0-shot prompts for each model. App.s D.1 and D.3 provide additional details.
@@
Metrics. For SEAM and ZeroScrolls we employ the metrics and corresponding implementations defined by each for their component datasets. For WikiHop and HotpotQA-distractor, we use the standard F1 score. For Multi-XScience and QMDSCNN, we utilize LLM scores issued by GPT-3.5-Turbo and Gemini 1.5 Flash, respectively, to assess summarization coherence, relevance, fluency, and consistency. This follows recent adoption of LLM-as-a-Judge as a proxy for human-level rating (Liu et al., 2024b; Ye et al., 2024), given the limitations of traditional metrics like ROUGE for long-form generation quality assessment. (Sai et al., 2019; Gao et al., 2024a). Further details regarding the evaluator LLMs and the prompts are in App. D.4.
@@ 5 Results
Model / IT Data Setting HQA WH MXS QC SEAM ZS Score Avg
@@
FlanT5 Base PRIMERA 0.4 0.5 70.7 24.2 7.1 7.7 0.5 3.9 10.4 5.9 4.2 8.8
QAMDen 1.8 1.9 63.6 27.1 0.0 0.4 0.0 1.0 3.1 0.9 3.8 7.2
None 4.4 45.1 38.7 48.0 0.0 5.8 0.2 2.6 0.1 1.7 13.1 14.5
Unfiltered 31.7 47.5 89.8 52.2 0.0 6.7 0.1 2.7 0.2 1.9 21.0 23.2
GPT-Filtered 44.1 46.0 92.4 54.2 0.0 6.0 0.2 2.7 0.0 1.8 21.3 24.1
MDCureRM 47.3 48.3 93.8 57.3 0.0 7.5 0.3 2.7 0.0 2.1 22.6 25.4
Large None 24.4 54.6 70.9 61.8 1.1 7.9 0.1 3.1 0.3 2.5 23.2 24.0
Unfiltered 46.9 62.9 91.1 64.7 0.7 8.0 0.0 3.8 0.1 2.5 24.2 27.3
GPT-Filtered 46.3 65.1 91.8 64.0 0.5 8.3 0.0 3.9 0.1 2.6 24.2 27.5
MDCureRM 49.6 66.1 93.1 66.0 1.2 9.1 0.1 4.2 0.0 2.9 25.3 28.5
Qwen2-Ins 1.5B None 21.5 17.8 93.3 73.3 9.4 10.6 0.5 5.0 13.0 7.7 24.0 25.5
Unfiltered 32.9 30.5 94.2 79.3 15.7 12.0 0.4 5.0 16.8 10.0 23.9 27.7
GPT-Filtered 33.3 32.9 94.2 81.3 16.1 12.0 0.5 5.1 15.9 9.9 24.2 28.1
MDCureRM 37.7 34.8 94.4 82.9 16.4 12.0 0.6 5.9 18.1 10.6 25.6 29.4
7B LongAlign-7B 10.4 14.3 92.2 83.3 11.5 16.5 0.0 4.1 16.8 9.8 23.2 25.3
None 30.5 39.6 95.6 79.3 5.0 11.9 0.5 6.4 13.1 7.4 23.9 27.4
Unfiltered 40.5 43.3 94.7 84.2 8.0 15.3 0.5 6.6 11.9 8.5 26.3 29.9
GPT-Filtered 42.0 44.0 94.7 85.3 8.7 15.3 0.9 6.6 12.1 8.7 28.7 31.4
MDCureRM 44.7 46.0 95.1 87.3 13.8 15.4 0.6 6.7 14.9 10.3 29.8 32.7
Llama3.1-Ins 8B ProLong-8B 43.6 34.3 85.8 54.7 9.5 17.4 0.4 10.7 18.0 11.2 32.2 32.1
None 35.5 27.1 95.1 65.3 10.5 15.0 0.6 7.6 17.4 10.2 18.7 24.3
Unfiltered 37.6 34.4 84.7 90.4 15.3 16.3 0.5 7.8 18.9 11.8 28.6 31.1
GPT-Filtered 38.0 42.3 95.3 87.8 8.7 16.0 0.6 7.8 18.3 10.3 29.6 32.1
MDCureRM 44.7 43.7 95.3 93.8 16.3 16.4 0.6 7.9 18.5 11.9 30.9 34.0
70B Jamba 1.5 Mini 47.5 41.8 94.2 87.1 20.1 14.2 0.0 5.3 20.4 12.0 34.1 35.3
None 53.9 38.1 95.1 88.2 25.1 21.9 0.6 6.1 11.5 13.0 36.4 37.1
Unfiltered 55.9 40.6 88.7 70.4 3.6 21.9 0.6 6.3 11.2 8.7 34.9 34.1
GPT-Filtered 57.4 41.2 88.2 74.9 4.9 22.0 0.7 6.4 10.7 8.9 37.5 35.9
MDCureRM 58.4 45.5 95.1 88.7 25.2 22.1 0.7 6.7 12.0 13.3 37.7 38.5
GPT-4o 57.5 50.0 100.0 94.4 8.9 18.6 0.3 14.2 28.9 14.1 39.8 40.6
Gemini 1.5 Pro 66.6 55.6 93.8 79.8 21.4 17.8 0.6 15.1 30.1 17.0 32.0 36.9
Table 1: Evaluation of MDCure’d models versus baselines on 6 benchmarks in the zero-shot prompting setting. The rightmost “Avg” column reports the average of individual dataset scores. Dataset abbreviations are described in App. D.1.1. Full results on ZeroScrolls are provided in App. F. Rows specified by “MDCureRM” refer to our full MDCure pipeline applied to the corresponding base model and size. Size-comparable baselines are highlighted in blue and results of our final method is highlighted in yellow. Bold numbers show best performance in each group.
@@ 5.1 Main Results
We report experimental results in Table 1, with full results on ZeroScrolls included in App. F. For fair comparison, we report all MDCure results based on training with 72K samples, which led to best performance overall across all models. Our key findings are summarized as follows.
MDCure instructions are effective across model families and sizes. MDCure exhibits the best performance across all base models, demonstrating a clear and consistent advantage relative to non-MDCure’d base models over all benchmarks. In particular, we see improvements of up to 975%, 96%, 143%, 434%, 39%, and 73% for HotpotQA, WikiHop, Multi-XScience, QMDSCNN, SEAM, and ZeroScrolls, respectively. MDCure’d LLAMA3.1-70B-Instruct achieves the best results across benchmarks, followed by MDCure’d Qwen2-7B-Instruct. As our models are trained from -instruct versions of base LLMs, it is clear that MDCure instruction data imparts complementary abilities versus typical IT data.
The relative improvement of MDCure’d models versus corresponding base models varies across model families, and improvements tend to shrink with model size. For FlanT5, MDCure provides 75.1% average improvement at the 250M-parameter scale (FlanT5-Base) versus 18.9% at the 750M-parameter scale (FlanT5-Large). For LLAMA3.1, MDCure provides 40.2% average improvement at the 8B scale versus 3.8% at the 70B scale. For Qwen2, we observe alternatively that MDCure gains shift from 15.4% on average at the 1.5B scale to 19.4% at the 7B scale. More generally, these results suggest the importance of instruction quality may weaken with base model scale and capability, similar to findings by Ivison et al. (2023).
Filtering with MDCureRM is essential to improve MD capabilities of LLMs. Comparing to models trained with unfiltered MDCure instruction samples, MDCure’d models trained with the highest-scoring MDCureRM-filtered instructions are superior across all benchmarks and base models. Furthermore, MDCureRM outperforms GPT-3.5-Turbo as a judge of MD data quality and efficacy, evidenced by MDCure’d models superiority over those trained with GPT-filtered instructions. As use of naive, unfiltered instructions yields relatively little improvement to MD capabilities especially at larger base model scales, this suggests curation of high-quality synthetic MD data is a nontrivial goal which MDCure is able to efficaciously accomplish.
MDCure improves performance on both MD and single-document long-context tasks, enabling cross-task and cross-domain generalization to unseen tasks and domains. Across MD benchmarks, MDCure enables models to achieve improved performance in all settings, with strong generalization to out-of-distribution (OOD) tasks such as MD coreference resolution, MD classification, text reordering, and aggregated sentiment classification and to OOD domains such as science, literature, and media. Beyond MD tasks, MDCure generally boosts single-document long-context performance, particularly on tasks in ZeroScrolls, which emphasize long-context abilities. These trends are consistent across downstream task domains, suggesting MDCure data boosts MD performance in a domain-general way.
@@
ProLong (PL) 43.6 34.3 85.8 54.7 11.2 33.6 32.1
PL+MDCure 45.8 45.1 87.3 75.6 12.5 34.7 34.9
Table 2: MDCure performance when applied to an already strong long-context LLM, ProLong-8B.
Data Source HQA WH MXS QC SEAM ZSS Avg
Qwen2-7B-Ins 30.5 39.6 95.6 79.3 7.4 23.9 27.4
+MDCure 44.7 46.0 95.1 87.3 10.3 29.8 32.7
+PRIMERA 44.6 45.4 88.0 48.2 8.1 27.9 28.7
+QAMDen 42.6 45.8 87.1 52.7 8.8 27.2 28.6
Table 3: Efficacy of MDCure data versus synthetic MD pre-training data, applied to Qwen2-7B-Instruct.
MDCure consistently outperforms pre-trained long-context baselines. Our MDCure’d models achieve results that exceed the performance of size-comparable baselines [^7] We consider Jamba 1.5 Mini (54B) comparable with LLAMA3.1-70B-Instruct, as their parameter scales are within the same ballpark, falling within the same order of magnitude. trained on much longer inputs (64-256K tokens) for all benchmarks by a substantial margin, confirming the efficacy of our design. For a controlled setup, we apply MDCure to strong open-source long-context LLM Prolong-8B (Gao et al., 2024a) to study if our data provides additional advantages to long-context continual pre-training and fine-tuning. Table 2 demonstrates the results. Note that Prolong-8B has undergone extensive training on $\sim$81B tokens to improve long-context performance. Fine-tuning ProLong-8B with much less MDCure data ($\sim$162M tokens) yields further improvement, verifying the added long-context value of our data.
Our MDCure pipeline improves over existing MD synthetic pre-training data. We also compare our MDCure pipeline with existing MD-focused synthetic pre-training datasets inluding PRIMERA and QAMDen. As indicated in Table 3, MDCure data offers performance improvements across benchmarks at the 7B-scale. The weak performance of the original PRIMERA and QAMDen models (Table 1) can be explained as they are intended to serve as strong base pre-trained models for finetuning as opposed to zero-shot prompting settings. Overall, our results emphasize the utility of our framework as a powerful approach for improving MD abilities of LLMs via post-training.